# Limiting Sets for Digital Cones and Suspensions

### v2

@@ 2.1 Adjacencies
A sequence $P=\{y_{i}\}_{i=0}^{m}$ in a digital image $(X,\kappa)$ is a $\kappa$-path from $a\in X$ to $b\in X$ if $a=y_{0}$, $b=y_{m}$, and $y_{i}\leftrightarroweq_{\kappa}y_{i+1}$ for $0\leq i<m$.
@@
A (digital) $\kappa$-closed curve is a path $S=\{s_{i}\}_{i=0}^{m}$ such that $s_{0}=s_{m}$, and $0<|i-j|<m$ implies $s_{i}\neq s_{j}$. We say $m$ is the length of the path:
$m=length(S).$
If, also, $0\leq i<m$ implies
$N(S,x_{i},\kappa)=\{x_{(i-1)\,mod\,m},~{}x_{(i+1)\,mod\,m}\}$
@@ 2.2 Digitally continuous functions
Digital continuity preserves connectedness, as at Definition 2.2 below. By using adjacency as our notion of “closeness,” we get Theorem 2.3 below.
@@ 2.4 Cones and suspensions
  • The cone on $X$ is the set $CX=X\cup\{U\}$, where $U\not\in X$, and adjacency $\kappa^{\prime}$, where $U\leftrightarrow_{\kappa^{\prime}}x$ for all $x\in X$; and if $a,b\in X$ then $a\leftrightarrow_{\kappa^{\prime}}b$ if and only if $a\leftrightarrow_{\kappa}b$.
@@ 3 Limiting sets for $CX$ and $SX$
Assume $D$ is a $\kappa$-dominating set in $X$. For each $x\in X$ there exists $\delta_{x}\in X$ such that $x\leftrightarroweq_{\kappa}\delta_{x}$. By continuity
@@
In $SX$, if $a\neq U$ then $\{a,L,b\}$ is a path in $SX$ from $a$ to $b$ of length at most 2. If $a=U$, then for any $w\in X$, $\{U,w,b\}$ is a path in $SX$ from $a$ to $b$ of length at most 2. ∎
@@
$f(x)=\left\{\begin{array}[]{ll}U&\mbox{if }x=x_{0};\\ x&\mbox{if }x\neq x_{0},\end{array}\right.$
@@
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\in X;\\ y&\mbox{if }x=U,\end{array}\right.$
@@
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq U;\\ L&\mbox{if }x=U\end{array}\right.$
@@
$g(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq L;\\ U&\mbox{if }x=L,\end{array}\right.$
is easily seen to be continuous, satisfies $g|_{B}=\operatorname{id}_{B}$, and $g\neq\operatorname{id}_{SX}$. Thus we must have $L\in B$. ∎
@@
Suppose $A$ is a minimal freezing set for $(X,\kappa)$. Let $B^{\prime}\subset B$ such that $B^{\prime}$ is a freezing set for $(SX,\kappa^{\prime})$. Let $A^{\prime}=B^{\prime}\setminus\{U,L\}\subset A$. Suppose $f\in C(X,\kappa)$ such that $f|_{A^{\prime}}=\operatorname{id}_{A^{\prime}}$. Let $g:CX\to CX$ be defined by
$g(y)=\left\{\begin{array}[]{ll}y&\mbox{if }y\in A^{\prime}\cup\{U,L\};\\ f(y)&\mbox{otherwise.}\end{array}\right.$
We see easily that $g\in C(SX,\kappa^{\prime})$ and $g|_{B^{\prime}}=\operatorname{id}_{B^{\prime}}$. By choice of $B^{\prime}$, $g=\operatorname{id}_{CX}$. Therefore, $f=\operatorname{id}_{X}$. Thus $A^{\prime}$ is a freezing set for $(X,\kappa)$. Since $A^{\prime}\subset A$ and $A$ is minimal, we have $A^{\prime}=A$, so by Proposition 3.6, $B^{\prime}=B$. Thus $B$ is a minimal freezing set for $(SX,\kappa^{\prime})$.
@@
$g(y)=\left\{\begin{array}[]{ll}f(y)&\mbox{if }y\in X;\\ y&\mbox{if }y\in\{U,L\},\end{array}\right.$
@@
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq U;\\ L&\mbox{if }x=U.\end{array}\right.$
@@ 4.1 Freezing sets for $P_{n}$
$f(p)=\left\{\begin{array}[]{ll}(n-1,-n+1,1)&\mbox{if }p=x;\\ p&\mbox{if }p\neq x\end{array}\right.$
@@
$f(p)=\left\{\begin{array}[]{ll}(n-1,z,1)&\mbox{if }p=x;\\ p&\mbox{if }p\neq x\end{array}\right.$
@@ 4.2 Freezing sets for digital solid pyramids
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y;\\ (0,0,n-1)&\mbox{if }x=U,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y;\\ (a,b,1)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y;\\ (-(n-1),b,0)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y;\\ (n-1,b,0)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y;\\ (a,-(n-1),0)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y;\\ (a,n-1,0)&\mbox{if }x=y,\end{array}\right.$
@@
For all $x\in Q_{n}$ there is a $c_{3}$-path $P=\{x_{n}\}_{n=0}^{n}$ from a point $x_{0}\in W_{n}$ to $U$ such that $x\in P$ and $p_{3}(x_{i+1})=p_{3}(x_{i})+1$ for $0\leq i<n$. It follows from Lemma 2.8 that if $p_{3}(f(x))<p_{3}(x)$, then pulling points of $P$ to lesser $3^{rd}$ coordinates would propagate so that $p_{3}(f(U))<p_{3}(U)$, contrary to $U\in\operatorname{Fix}(f)$; and if $p_{3}(f(x))>p_{3}(x)$, then pulling points of $P$ to greater $3^{rd}$ coordinates would propagate so that $p_{3}(f(x_{0}))>p_{3}(x_{0})$, contrary to $x_{0}\in W_{n}\subset\operatorname{Fix}(f)$. Thus, (8) is established.
We show $f$ satisfies
$p_{1}(f(x))=p_{1}(x)\mbox{ for all }x\in Q_{n}.$ (9)
Let $x=(a,b,c)\in Q_{n}$. Then $x\in\{\{(j,b,c)\}_{j=-(n-c)}^{n-c}\}$, a digital line segment from $(-(n-c),b,c)\in B\subset\operatorname{Fix}(f)$ to $(n-c,b,c)\in F\subset\operatorname{Fix}(f)$. As above, from Lemma 2.8, we must have $x$ satisfying (9).
A similar argument shows
$p_{2}(f(x))=p_{2}(x)\mbox{ for all }x\in Q_{n}.$ (10)
From (8), (9), and (10), we have $Q_{n}\subset\operatorname{Fix}(f)$, so $f=\operatorname{id}_{Q_{n}}$. Hence $A$ is freezing. ∎
@@ 5 Freezing sets for suspension-like images
We show $A$ is freezing, as follows. Let $f\in C(H_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$.
@@
Similarly, $f|_{P_{n}^{\prime}}=\operatorname{id}_{P_{n}^{\prime}}$. It follows that $A$ is a freezing set for $(H_{n},c_{3})$. ∎
@@
Let $A$ be a freezing set for $(K_{n},c_{3})$. If $U\not\in A$ then the function $f:K_{n}\to K_{n}$ given by
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq U;\\ (0,0,n-1)&\mbox{if }x=U,\end{array}\right.$
is easily seen to belong to $C(K_{n},c_{3})$, with $f|_{A}=\operatorname{id}_{A}$ and $f\neq\operatorname{id}_{K_{n}}$. The contradiction implies $U\in A$.
Similarly, $L\in A$.
Since $T_{n}=BL\cup BF\cup BR\cup BB$, we can show $T_{n}\subset A$ by using the argument given in the proof of Proposition 4.3. ∎
@@
Given $x\in K_{n}\setminus\{U,L\}$, if $p_{3}(f(x))\neq p_{3}(x)$ then either
$d(x,U)<d(f(x),U)=d(f(x),f(U)),$
or
$d(x,L)<d(f(x),L)=d(f(x),f(L)).$
Both are impossible, since $f$ is continuous. Therefore

### v3

@@ 2.1 Adjacencies
A sequence $P=\{y_{i}\}_{i=0}^{m}$ in a digital image $(X,\kappa)$ is a $\kappa$-path from $a\in X$ to $b\in X$ if $a=y_{0}$, $b=y_{m}$, and $y_{i}\leftrightarroweq_{\kappa}y_{i+1}$ for $0\leq i<m$. We say $m$ is the length of the path:
$m=length(S).$
@@
A (digital) $\kappa$-closed curve is a path $S=\{s_{i}\}_{i=0}^{m}$ such that $s_{0}=s_{m}$, and $0<|i-j|<m$ implies $s_{i}\neq s_{j}$. If, also, $0\leq i<m$ implies
$N(S,x_{i},\kappa)\setminus\{x_{i}\}=\{x_{(i-1)\,mod\,m},~{}x_{(i+1)\,mod\,m}\}$
@@ 2.2 Digitally continuous functions
Digital continuity preserves connectedness, as at Definition 2.2. By using adjacency as our notion of “closeness,” we get Theorem 2.3.
@@ 2.4 Cones and suspensions
  • The cone on $X$ is the set $CX=X\cup\{U\}$, where $U\not\in X$, and adjacency $\kappa^{\prime}$, where $U\leftrightarrow_{\kappa^{\prime}}x$ for all $x\in X$ and if $a,b\in X$ then $a\leftrightarrow_{\kappa^{\prime}}b$ if and only if $a\leftrightarrow_{\kappa}b$.
@@ 3 Limiting sets for $CX$ and $SX$
Assume $D$ is a $\kappa$-dominating set in $X$. For each $x\in X$, there exists $\delta_{x}\in D$ such that $x\leftrightarroweq_{\kappa}\delta_{x}$. By continuity
@@
In $SX$, for $\{a,b\}\subset X$, $\{a,U,b\}$ is a path in $SX$ from $a$ to $b$ of length 2. For $a\in X$, $\{a,L\}$ is a path in $SX$ of length 1 from $a$ to $L$. For $b\in X$, $\{U,b\}$ is a path in $SX$ of length 1 from $U$ to $b$. For any $x\in X$, $\{U,x,L\}$ is a path in $SX$ from $U$ to $L$ of length 2. The assertion follows. ∎
@@
$f(x)=\left\{\begin{array}[]{ll}U&\mbox{if }x=x_{0},\\ x&\mbox{if }x\neq x_{0},\end{array}\right.$
@@
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\in X,\\ y&\mbox{if }x=U,\end{array}\right.$
@@
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq U,\\ L&\mbox{if }x=U\end{array}\right.$
@@
$g(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq L,\\ U&\mbox{if }x=L,\end{array}\right.$
is easily seen to be a continuous function that satisfies $g|_{B}=\operatorname{id}_{B}$, and $g\neq\operatorname{id}_{SX}$. Thus we must have $L\in B$. ∎
@@
Suppose $A$ is a minimal freezing set for $(X,\kappa)$. Let $B^{\prime}\subset B$ such that $B^{\prime}$ is a freezing set for $(SX,\kappa^{\prime})$. Let $A^{\prime}=B^{\prime}\setminus\{U,L\}\subset A$. Suppose $f\in C(X,\kappa)$ such that $f|_{A^{\prime}}=\operatorname{id}_{A^{\prime}}$. Let $g:SX\to SX$ be defined by
$g(y)=\left\{\begin{array}[]{ll}y&\mbox{if }y\in A^{\prime}\cup\{U,L\},\\ f(y)&\mbox{otherwise.}\end{array}\right.$
We see easily that $g\in C(SX,\kappa^{\prime})$ and $g|_{B^{\prime}}=\operatorname{id}_{B^{\prime}}$. By choice of $B^{\prime}$, $g=\operatorname{id}_{SX}$. Therefore, $f=\operatorname{id}_{X}$. Thus $A^{\prime}$ is a freezing set for $(X,\kappa)$. Since $A^{\prime}\subset A$ and $A$ is minimal, we have $A^{\prime}=A$, so by Proposition 3.6, $B^{\prime}=B$. Thus $B$ is a minimal freezing set for $(SX,\kappa^{\prime})$.
@@
$g(y)=\left\{\begin{array}[]{ll}f(y)&\mbox{if }y\in X,\\ y&\mbox{if }y\in\{U,L\},\end{array}\right.$
@@
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq U,\\ L&\mbox{if }x=U.\end{array}\right.$
@@ 4.1 Freezing sets for $P_{n}$
$f(p)=\left\{\begin{array}[]{ll}(n-1,-n+1,1)&\mbox{if }p=x,\\ p&\mbox{if }p\neq x\end{array}\right.$
@@
$f(p)=\left\{\begin{array}[]{ll}(n-1,z,1)&\mbox{if }p=x,\\ p&\mbox{if }p\neq x\end{array}\right.$
@@ 4.2 Freezing sets for digital solid pyramids
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (0,0,n-1)&\mbox{if }x=U,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (a,b,1)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (-(n-1),b,0)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (n-1,b,0)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (a,-(n-1),0)&\mbox{if }x=y,\end{array}\right.$
@@
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (a,n-1,0)&\mbox{if }x=y,\end{array}\right.$
@@
For all $x\in Q_{n}$ there is a $c_{3}$-path $P=\{x_{i}\}_{i=0}^{n}$ from a point $x_{0}\in W_{n}$ to $U$ such that $x\in P$ and $p_{3}(x_{i})=i$ for $0\leq i\leq n$. It follows from Lemma 2.8 that if $p_{3}(f(x))<p_{3}(x)$, then pulling points of $P$ to lesser $3^{rd}$ coordinates would propagate so that $p_{3}(f(U))<p_{3}(U)$, contrary to $U\in\operatorname{Fix}(f)$; and if $p_{3}(f(x))>p_{3}(x)$, then pulling points of $P$ to greater $3^{rd}$ coordinates would propagate so that $p_{3}(f(x_{0}))>p_{3}(x_{0})$, contrary to $x_{0}\in W_{n}\subset\operatorname{Fix}(f)$. Thus, (8) is established.
Next, we show
$LR\cup LF\cup RF\cup RR\subset\operatorname{Fix}(f).$ (9)
To establish the latter claim, we will show $LF\subset\operatorname{Fix}(f)$. Notice $LF$ is the unique shortest $c_{3}$-path in $Q_{n}$ from $(n,-n,0)\in A$ to $U\in A$. By Proposition 2.7, $LF\subset\operatorname{Fix}(f)$ A similar argument shows that $LR\cup RF\cup RR\subset\operatorname{Fix}(f)$, and (9) follows.
Now, we claim
$\begin{array}[]{l}u\in L\cup R\mbox{ implies }p_{1}(f(u))=p_{1}(u);\\ v\in F\cup B\mbox{ implies }p_{2}(f(v))=p_{2}(v).\end{array}$ (10)
Suppose $u=(a,b,c)\in L$. Then $u\in\{(i,b,c)\}_{i=-(n-c)}^{(n-c)}$, a $c_{3}$-path from $u_{0}=(-(n-c),b,n-c)\in LR\subset\operatorname{Fix}(f)$ to $u^{\prime}=(n-c,b,n-c)\in LF\subset\operatorname{Fix}(f)$. If $p_{1}(f(u))>p_{1}(u)$ then by Lemma 2.8 we would have $p_{1}(f(u_{0}))>p_{1}(u_{0})$, contrary to $u_{0}\in\operatorname{Fix}(f)$. If $p_{1}(f(u))<p_{1}(u)$ then by Lemma 2.8 we would have $p_{1}(f(u^{\prime}))<p_{1}(u^{\prime})$, contrary to $u^{\prime}\in\operatorname{Fix}(f)$. Hence $p_{1}(f(u))=p_{1}(u)$. We show the other claims of (10) similarly.
Now let $q=(a,b,c)\in Q_{n}$. Then $q$ belongs to the paths
$P_{1}=\{(i,b,c)\}_{i=-(n-c)}^{n-c},~{}~{}~{}~{}~{}P_{2}=\{(a,j,c)\}_{j=-(n-c)}^{n-c}$
with endpoints in $B\cup F$ and in $L\cup R$, respectively. Note the sequence of points in $P_{1}$ is monotone increasing in the first coordinate. It follows as above from (10) and Lemma 2.8 that we must have $p_{1}(f(q))=p_{1}(q)$. Similarly, $p_{2}(f(q))=p_{2}(q)$.
From (8), it follows that $f(q)=q$, i.e., $f=\operatorname{id}_{Q_{n}}$. Hence $A$ is freezing. ∎
@@ 5 Freezing sets for suspension-like images
We show $T_{n}$ is freezing, as follows. Let $f\in C(H_{n},c_{3})$ such that $f|_{T_{n}}=\operatorname{id}_{T_{n}}$.
@@
Similarly, $f|_{P_{n}^{\prime}}=\operatorname{id}_{P_{n}^{\prime}}$. It follows that $T_{n}$ is a freezing set for $(H_{n},c_{3})$. ∎
@@
The assertion follows as in the proof of Proposition 4.3. ∎
@@
Given $x\in K_{n}\setminus\{U,L\}$, we have as in the proof of Theorem 4.4 that